2. A man who weighs 600 N is sitting in a chair with his feet on the floor and arms resting on the armrests. The chair weighs 100 N. Each armrest exerts and upward force of 25.0 N on his arms, and the seat exerts an upward force of 500 N. What force does the floor exert on his feet?
1
Expert's answer
2019-05-13T10:47:32-0400

From third Newton's law:


"W=F_{upward}"

"F_{upward}=F_1+F_2+2F_3"

"F_1=W-F_2-2F_3=600-500-2(25)=50\ N"